If a an b are two positive numbers such that a+b=640,where a is 100 greater than b,then a and b ?

Grade:8

1 Answers

tanmay pratap singh
6 years ago
We know that a=100+b anda+b=640 by evaluting the value of a=100+b in this equation we get 100+b+b=640So b=270 and a=100+270=370